Calculate the AH formation of octane, CBH1B (), given that the enthalpy of combustion of octane is -5471 kl/mol and the standard enthalpies of formation of H20 () and CO2 (g) are -393.5 kJ/mol and -285.8 kJ/mol respectively.
Answer #1

Chemical reaction for combustion of C8H18 is ;

2C8H18 + 25O2 = 18H2O + 16CO2 + heat (given -5471kJ/mol)

?H of reaction = sum of ?H°f of formation of products - sum of ?H°f formation of reactants .

2× -5471(as 2 moles ) = ( ?H°f 18 H2O + ?H°f 16 CO2 ) - ?H f of 2 Octane ( ?H°f of O2 is zero) .

?H°f of 2 Octane = [18 (-393.5) + 16 (-285.8)] - (-10942)

?H°f of 2 moles of Octane = - 11655.8 + 10942 = -713.8 kJ .

Now enthalpy of formation of octane(l) = -713.8/2 = -356.9 kJ/mole .